Consider two bonds. The first is a 6% coupon bond with six years to maturity, and a yield to maturity of 4.5% annual rate, compounded semi-annually. The second bond is a 2% coupon bond with six years to maturity and a yield to maturity of 5.0%, annual rate, compounded semi-annually.

1. Calculate the current price per $100 of face value of each bond. (You may use financial calculator to do question 1 and 2, I'm just unsure how to use it.)

2. Given the data for the first two bonds, now consider a third bond: a zero coupon bond with six years to maturity. Calculate the price per $100 of face value of the zero coupon bond. Calculate the yield to maturity for the zero coupon bond. (Express the yield as annual rate, compounded semi-annually).

HINT: Use the Value Additivity principle to answer part c. Create a synthetic zero- coupon bond, that is, a portfolio of the 6% coupon bond and the 2% coupon bond that has the same cash flows as a 6-year, zero coupon bond.
Answer #1

For both the bonds, coupon is compounded semiannualy

a) First bond

Assuming Face value = 100, Semiannual coupon = 6% / 2* 100 =3

Yield to maturity = 4.5%

Years to maturity = 6 , Number of coupon payment period = 6 * 2 = 12

Using excel, Bond current price = $86.3

Period Coupon Yield / discount rate Discount factor Present value of cash flow
1 3 4.50% 0.957 2.9
2 3 4.50% 0.916 2.7
3 3 4.50% 0.876 2.6
4 3 4.50% 0.839 2.5
5 3 4.50% 0.802 2.4
6 3 4.50% 0.768 2.3
7 3 4.50% 0.735 2.2
8 3 4.50% 0.703 2.1
9 3 4.50% 0.673 2.0
10 3 4.50% 0.644 1.9
11 3 4.50% 0.616 1.8
12 103 4.50% 0.590 60.7
Bond present value 86.3

Excel formula:

B A C D E 19 Period Coupon Yield/discount rate Discount factor Present value of cash flow - 100 6 % / 2 20 1 =1/(1+C20)^A20 -

b) For second bond coupon is 2% (compounded semmiannually) and Yield to maturity is 5%, years to maturity =6%

Using excel, Bond current price = $64.5

Period Coupon Yield / discount rate Discount factor Present value of cash flow
1 1 5.00% 0.952 1.0
2 1 5.00% 0.907 0.9
3 1 5.00% 0.864 0.9
4 1 5.00% 0.823 0.8
5 1 5.00% 0.784 0.8
6 1 5.00% 0.746 0.7
7 1 5.00% 0.711 0.7
8 1 5.00% 0.677 0.7
9 1 5.00% 0.645 0.6
10 1 5.00% 0.614 0.6
11 1 5.00% 0.585 0.6
12 101 5.00% 0.557 56.2
Bond present value 64.5

2) Using addition principal

Price of zero coupon bond = average (value of bond 1, value of bond 2)

= average(86.3,64.5)

Price of zero coupon bond = $75.4

Using rate formula to to calculate yield = RATE(6,0,-75.4,100)

Yield on zero coupon = 4.8%